Evaluate the function y=1/2x-4 for each of the given domain values.
1. f(-2)=
2. f(0)=
3. f(3/2)=
4. f(-2/3)=
How do you answer these types of questions? Please help me with this!

Answer by rothauserc(4718)

f(x) = y = 1/2x-4
:
the order of operations(multiplication and division first from left to right) tells us that
:
f(x) = (1/2) * x -4
:
now just substitute the values for x and evaluate f(x)
:
1. f(-2) = (1/2) * -2 -4 = -5
:
2. f(0) = (1/2) * 0 -4 = -4
:
3. f(3/2) = (1/2) * (3/2) -4 = -13/4
:
4. f(-2/3) = (1/2) * (-2/3) -4 = -26/6 = -13/3
: